OpenStudy (anonymous):

log (base 7) 24701 =5x+8 the solution set is ______

OpenStudy (anonymous):

is it your new?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to me, i have 8 = log (base 7) of 5764801

OpenStudy (anonymous):

and move it into left side to have 5x = log(base 7) - new log . hey kid , do it next

OpenStudy (anonymous):

@Hoa you are doing just great.. I am an audience

OpenStudy (anonymous):

i don't think you cannot step up from then. do it. if you don't know how to calculate on calculator, let everything there. it's ok for the result. I mean (log - log )/5 =x

OpenStudy (anonymous):

@flutterflies are you with me? if you lost, just tell me . I will redo.

OpenStudy (anonymous):

how did you get 5764801?

OpenStudy (anonymous):

oh, since log(base7) of something = 8 so, 7^8 = that (something) . i just want to make the same log 7. if you don't want it. you can NOT to do it, just let log(7) of 24... -8 =5x . it's ok , too

OpenStudy (anonymous):

okay so then how would you put it in a calculator?

OpenStudy (anonymous):

I did. but it's not an integer, so you are better let it as it is

OpenStudy (anonymous):

ok, I put them in neat for you

OpenStudy (anonymous):

\[\log_{7} 24701= 5x +8 ---> 5x = \log_{7} 24701-8-->x=\frac{ \log_{7} 24701-8 }{ 5 }\]

OpenStudy (anonymous):

so that s the solution. and it cant get any more simplified. thank you for taking your time to show me! and hahaha yes i am a girl and that is me. Thank you!!

OpenStudy (anonymous):

is there a chance you can help me with just three more problems?

OpenStudy (anonymous):

I have a friend who is more intelligent than me, just post it. let me check whether he is online or not, if he isn't , I will